OpenStudy (anonymous):

The half-life of Radium-226 is 1590 years. If a sample contains 300 mg, how many mg will remain after 1000 years

OpenStudy (anonymous):

\[300\times \left(\frac{1}{2}\right)^{\frac{1000}{1590}}\] and a calculator